OpenStudy (anonymous):

Solving Two Step Equations/Show work/Check 6m+1=-23

OpenStudy (anonymous):

as with before 6m+1=23 6m=22 m=22/6 CHECK-------------------- 6(22/6)+1=23 22+1=23 23=23 the 6 * 22/6 cancel the 6 so 22.

OpenStudy (anonymous):

If -23 - -1 wouldn't that be -24?

OpenStudy (anonymous):

Where did you get -23?

OpenStudy (anonymous):

The answer is -23.